OpenStudy (mrnebula):

Find the simplified form of the expression. Give your answer in scientific notation. (3 * 10^6)(8 * 10^-4) the answer I got is (3 x 10^6) x 8/10^4 but that's not any of the answers...

OpenStudy (anonymous):

Because that's not the answer.

OpenStudy (anonymous):

You have to give your answer in scientific notation.

OpenStudy (mrnebula):

explain

OpenStudy (phi):

(3 * 10^6)(8 * 10^-4) this is the same as multiplying \[ 3 \cdot 10^6 \cdot 8 \cdot 10^{-4} \] or \[ (3 \cdot 8) \cdot 10^6 \cdot 10^{-4} \] 3*8 is 24 multiplying the other two terms by adding the exponents the last step is divide 24 by 10 and multiply the other term by 10 so that the answer is in standard from